OpenStudy (anonymous):

True or False? Y varies directly with x and y = -9 when x = 3. When y = 27, the value of x is 9. True False

OpenStudy (johnweldon1993):

well if this is the case the equation would be y = -3x -9 = -3(3), -9 = -9 yes....but 27 = -3(9)...27 = -27? Noooo False

OpenStudy (anonymous):

thanks

OpenStudy (johnweldon1993):

no problem bud